OpenStudy (anonymous):

Given the equation y − 4 = (x + 8) in point-slope form, identify the equation of the same line in slope-intercept form. −x + y = 10 3x − 4y = −40 y = x + 12 y = x + 10

OpenStudy (anonymous):

\[y - 4 = x + 8\\y = x + 12 \]

OpenStudy (precal):

add 4 to both sides

OpenStudy (anonymous):

thx guys